I want to score 80% in a class, the first 3 tests scores are 71, 82, and 90. What is the lowest score I can get on the 4th test? I need to know how to set this problem up to solve.
Answer by rfer(16322)

(71+82+90+x)/4=80
243+x=80*4
x=320-243
x=77%